OpenStudy (anonymous):

HOW DO I SOLVE 6 + (pi)/2 cos [(pi)x/4]

OpenStudy (anonymous):

Is this equal to zero?

OpenStudy (anonymous):

yes the answer is (pi)+12/2 and 12-(p)/2

OpenStudy (anonymous):

I dont know how to get to the answer though

myininaya (myininaya):

\[6+\frac{\pi}{2}\cos(\frac{\pi x}{4})=0\]

myininaya (myininaya):

is the above written correctly? \[\frac{\pi}{2} \cos(\frac{ \pi x}{4})=-6\] \[\cos( \frac{\pi x}{4})=\frac{-12}{\pi}\]

OpenStudy (anonymous):

YES THANK YOU I just realised that all i need to do is |amplitude|+vertical translation